1) Suppose the following for a resource (measure in tons):

  • Market demand schedule is Pd = 200 - 4Q
  • Market supply schedule is Ps = 80 + 2Q.
  • Total Reserve is 30 tons

The equilibrium quantity will be __ and the market price will be __.

2)

Suppose (as in question 17) the following for a resource (measure in tons):

  • Market demand schedule is Pd = 200 - 4Q
  • Market supply schedule is Ps = 80 + 2Q.
  • Total Reserve is 30 tons

and

  • There are 2 periods (each period is 5 years)
  • The Discount rate is 6.5%

What is the socially efficient quantity for period 1?
Answer #1

At equilibrium demand equals with supply.
Pd = Ps
So, 200-4Q = 80 + 2Q
So, 120 = 6Q
So, Q = 20

Now putting the value of Q in supply equation now putting the value of Q in supply equation,

It is given that, Ps = 80 - 2Q
So, Ps = 80 - 2*20.
So, Ps = 80 - 40 = 40

So at Equilibrium
Quantity = 20
Price = 40

NB: For second part question number 17 is missing please provide that.